Please help me solve this word problem using the Pythagorean Theorem: Dirk rowed a boat across a river 225 feet wide. A strong current carried his boat 400 feet downstream as he rowed. Find x, the distance Dirk actually rowed to the nearest foot.
I tried doing: a^2 + b^2 = c^2 and ended up getting sqrt of 306 as my answer.. which isn't right.

Answer by nerdybill(7384)

Dirk rowed a boat across a river 225 feet wide. A strong current carried his boat 400 feet downstream as he rowed. Find x, the distance Dirk actually rowed to the nearest foot.
.
Let x = distance (feet) actually rowed
then
x^2 = 225^2 + 400^2
x^2 = 50625 + 160000
x^2 = 210625
x = 459 feet